# Instruct-ReID++: Towards Universal Purpose Instruction-Guided Person Re-identification

### v1

@@ Abstract
Human intelligence can retrieve any person according to both visual and language descriptions. However, the current computer vision community studies specific person re-identification (ReID) tasks in different scenarios separately, which limits the applications in the real world. This paper strives to resolve this problem by proposing a novel instruct-ReID task that requires the model to retrieve images according to the given image or language instructions. Instruct-ReID is the first exploration of a general ReID setting, where existing 6 ReID tasks can be viewed as special cases by assigning different instructions. To facilitate research in this new instruct-ReID task, we propose a large-scale OmniReID++ benchmark equipped with diverse data and comprehensive evaluation methods e.g., task-specific and task-free evaluation settings. In the task-specific evaluation setting, gallery sets are categorized according to specific ReID tasks. We propose a novel baseline model, IRM, with an adaptive triplet loss to handle various retrieval tasks within a unified framework. For task-free evaluation setting, where target person images are retrieved from task-agnostic gallery sets, we further propose a new method called IRM++ with novel memory bank-assisted learning. Extensive evaluations of IRM and IRM++ on OmniReID++ benchmark demonstrate the superiority of our proposed methods, achieving state-of-the-art performance on 10 test sets. The datasets, the model, and the code will be available at https://github.com/hwz-zju/Instruct-ReID.
@@
Task-free evaluation setting. When evaluating the generalization ability of the model across multiple retrieval tasks jointly, task-specific instructions for each gallery image are discarded. Therefore, we propose a novel task-free evaluation setting as an extension of our previous conference paper. As shown in Fig 2 (b), the test set maintains a consistent gallery input for all tasks. Each image in the gallery set is task-agnostic and only the image feature is extracted for various ReID tasks. The gallery features can be fixed and extracted only once in applications, thus saving more computational resources. -
@@
To address the instruct-ReID task, we present an Instruct-ReID Model (IRM) in our previous conference version that can, for the first time, handle the major 6 ReID tasks listed in Fig 1 (a) with a dual-branch framework. Unlike the typical triplet loss [25] that only defines positive/negative pairs by identities, our IRM incorporates a novel adaptive triplet loss to learn a metric space that preserves identity and instruction similarities. Specifically, we design an adaptive margin between two query-instruction pairs based on instruction similarities to pull features with similar instructions close and push features with different instructions apart. Although the previous IRM has achieved significant success in unifying the multiple ReID tasks within one framework, it can not be applied in the newly proposed task-free evaluation setting for the following two reasons. First, the gallery features of various ReID tasks are extracted from different modules of IRM, e.g., Trad-ReID from the output of the attention module and T2I-ReID from the image encoder. However, in the task-free evaluation setting, all images in the gallery set are task-agnostic, making it impractical to extract retrieval features from different model parts based on the task type of gallery images. Second, IRM extracts gallery images and specific instruction features for retrieval comparison, while these task-specific instructions are unavailable in the task-free evaluation setting. To address the challenges, we introduce a new method called IRM++, which employs simple and unified gallery features for various ReID tasks. Concretely, unlike IRM, IRM++ obtains the gallery features by only extracting the image features from diverse tasks using an image encoder, without introducing task-specific multimodality instruction inputs. Furthermore, IRM++ introduces two memory banks in a contrastive learning manner to effectively supervise the model learning. Unlike triplet loss-based methods that include only one negative sample per triplet, memory bank contrastive learning offers multiple negative samples for each iteration, enhancing the discriminative representation learning of the data.
@@ II-B Loss Function in ReID
Mulitple loss function have been extensively studied during the development of the ReID research [28]. Depending on the structure of the method design, there are three typical types of loss function: (1) Identity loss proposed in [30, 31, 32, 33, 34, 35] treats each identity as a distinct class, thus turning the training phase of ReID as a classification process. Identity loss using cross-entropy loss to compute the difference between the labels and the predicted probability. It’s convenient for training but easy to overfit when the labels are over-confident annotated [36]. (2) Contrastive loss [37, 38, 39, 40] has been widely used in diverse ReID scenarios. It utilizes Euclidean distance to compute the pairwise distance of two input sample features. Contrastive loss encourages features of the same identities to come closer together while pushing features from different identities farther apart. The Contrastive loss can alleviate the impact of wrong classifications and noise during the clustering stage. Although contrastive loss can be sensitive to the choice of positive and negative samples during training, it faces the challenge of lacking a reference distance between samples in ReID task applications. (3) Triplet loss [41, 42, 43, 44] is also an extensively explored feature clustering method. It uses a retrieval ranking perspective to address the training process of the ReID model [28]. Triplet loss takes one anchor sample, one positive sample, and one negative sample as input; based on the assumption that the distance between the anchor and the positive sample should be smaller than the distance between the anchor and the negative sample by a predefined margin, triplet loss optimizes the distance within the triplet samples during the training phase.
@@ III OmniReID++ Benchmark
| dataset | image | ID | domain |
| MSMT17 [3] | 30,248 | 1,041 | indoor/outdoor |
| Market1501 [23] | 12,936 | 751 | outdoor |
| PRCC [24] | 17,896 | 150 | indoor |
| COCAS+ Real1 [12] | 34,469 | 2,800 | indoor/outdoor |
| LLCM [62] | 30,921 | 713 | indoor/outdoor |
| LaST [63] | 71,248 | 5,000 | indoor/outdoor |
| MALS [61] | 1,510,330 | 1,510,330 | synthesis |
| LUPerson-T [64] | 957,606 | - | indoor/outdoor |
| OmniReID++ | 5,072,218 | 333,825 | indoor/outdoor/synthetic |
@@ V-C Task-specific Evaluation Setting Results
V-D Task-free Evaluation Setting Results
Single Test Set Evaluation. We evaluate the performance of IRM++ on each single test set, which is a subset of the task-free evaluation setting. As shown in Tab. VI, for clothes template based clothes-changing ReID (CTCC-ReID), language-instructed ReID (LI-ReID), clothes-changing ReID (CC-ReID), visible-infrared ReID (VI-ReID), text-to-image ReID (T2I-ReID), and traditional ReID (Trad-ReID), under Multi-task learning (MTL) training, our IRM++ achieves state-of-the-art (SOTA) results, surpassing the best-performing methods by +0.3% to +22.4% mAP. Experimental results demonstrate that IRM++ effectively extracts generic gallery features and compares them with features extracted from query images and instructions for accurate retrieval. Besides, we conduct a baseline method IRM++* by employing a triplet loss as the supervision function. For CTCC-ReID and LI-ReID tasks, IRM++ surpasses the baseline method by +10.0% and +9.0% mAP (STL), +7.1% and +5.8% mAP (MTL), which indicates that the memory bank assisted contrastive learning utilized in IRM++ further enhances the model performance compared to optimization with triplet loss on these two ReID tasks. We analyze that the memory bank-assisted contrastive loss employs multiple negative samples for each iteration, providing stronger supervision during training compared to triplet-based optimization methods, which only utilize one negative sample per iteration. For CC-ReID, VI-ReID, and Trad-ReID, our IRM++ and baseline method achieve similar results, with experimental variations within 0.3% mAP under Single-task learning (STL) training.
Joint Test Set Evaluation. As shown in Tab. VII, we unify all testing set data into a joint set and provide benchmark testing results on the proposed dataset. We also perform inference with existing state-of-the-art methods on this testing set for comparison. Since DualBCT Net [12], RaSa [18], PASS [26], CAL [7] and DEEN [62] cannot receive multimodal instruction inputs, e.g., clothes template images and description texts, and the instruction similarity is unavailable, we only measure the traditional mAP metric on the Joint Test dataset. We also provide the inference performance of the baseline method IRM++* on the newly proposed testing set. Under the traditional mAP evaluation metric, our IRM++ achieves the best result of 65.3% mAP, surpassing the baseline method and existing methods by 9.1% mAP and 22.1% mAP, respectively. The results demonstrate that IRM++ exhibits outstanding performance in the task-free evaluation setting. Based on instruction similarity and our proposed mAP evaluation metric, we provide evaluation results at three different thresholds of 0.25, 0.50, and 0.75. Compared to the baseline method, IRM++ achieves better results of +8.9% (mAP25), +6.7% (mAP50), and +6.3% (mAP75). We observe that under the mAP evaluation metric, the test results of our proposed IRM++ show a decreasing trend as the threshold increases. This indicates that with increasing thresholds, more gallery samples with the same identity as the query image but not matching the instruction attributes are corrected back to negative samples, aligning with our proposed mAP evaluation metric. The selection of thresholds for different models and scenarios is an area worthy of future research.
@@ V-E Ablation Study
| IRM | CTCC-ReID | LI-ReID | T2I-ReID | VI-ReID | CC-ReID | Trad-ReID | ||||||
| SOTA | - | 30.0 [12] | 14.9 | 63.9 [18] | 65.8 [62] | 62.9 [62] | 40.8 [7] | 45.9 [78] | 71.8 [79] | 93.0 [26] | 71.8 [26] | 77.7 [85] |
| DeiT [87] | STL(256128) | 31.5 | 30.1 | 64.2 | 65.1 | 62.3 | 46.3 | 43.1 | 82.1 | 87.9 | 67.9 | 77.5 |
| MTL(256128) | 40.2 | 38.7 | 65.7 | 66.2 | 62.9 | 53.2 | 46.8 | 76.3 | 90.0 | 68.9 | 78.7 | |
| ALBEF [74] | STL(256128) | 30.6 | 28.7 | 67.7 | 61.2 | 58.8 | 40.2 | 42.1 | 67.1 | 78.9 | 60.3 | 69.2 |
| MTL(256128) | 37.2 | 34.5 | 68.3 | 62.1 | 60.1 | 41.2 | 44.1 | 60.5 | 79.7 | 62.1 | 69.9 | |
| HAP [88] | STL(256128) | 30.8 | 31.1 | 63.2 | 66.2 | 62.9 | 45.3 | 44.7 | 79.2 | 90.2 | 70.3 | 79.6 |
| MTL(256128) | 39.2 | 39.8 | 65.1 | 67.1 | 64.3 | 49.2 | 48.7 | 77.3 | 91.7 | 72.3 | 82.2 | |
| PASS [26] | STL(256128) | 32.2 | 30.7 | 65.3 | 66.6 | 64.5 | 46.7 | 46.0 | 80.1 | 92.3 | 71.9 | 83.3 |
| MTL(256128) | 41.7 | 39.8 | 66.5 | 68.5 | 67.2 | 52.0 | 52.3 | 78.9 | 93.5 | 72.4 | 85.4 | |
| STL(384192) | 33.7 | 33.8 | 65.9 | 68.9 | 64.7 | 48.9 | 49.3 | 82.5 | 92.7 | 73.7 | 84.1 | |
| MTL(384192) | 42.5 | 42.3 | 67.2 | 71.2 | 67.3 | 54.1 | 58.4 | 81.8 | 93.9 | 75.4 | 88.7 | |
| IRM++ | CTCC-ReID | LI-ReID | T2I-ReID | VI-ReID | CC-ReID | Trad-ReID | ||||||
| SOTA | - | 30.0 [12] | 14.9 | 63.9 [18] | 65.8 [62] | 62.9 [62] | 40.8 [7] | 45.9 [78] | 71.8 [79] | 93.0 [26] | 71.8 [26] | 77.7 [85] |
| DeiT [87] | STL(256128) | 29.3 | 27.9 | 63.1 | 64.4 | 62.1 | 46.1 | 42.9 | 80.1 | 87.2 | 67.7 | 77.4 |
| MTL(256128) | 32.8 | 34.7 | 66.9 | 66.1 | 62.3 | 52.8 | 45.9 | 75.3 | 89.8 | 68.1 | 78.2 | |
| ALBEF [74] | STL(256128) | 27.6 | 25.4 | 66.9 | 61.1 | 58.7 | 40.1 | 42.2 | 66.7 | 78.7 | 60.1 | 68.8 |
| MTL(256128) | 29.9 | 30.2 | 68.1 | 63.0 | 60.3 | 41.1 | 44.1 | 59.2 | 79.3 | 62.0 | 69.4 | |
| HAP [88] | STL(256128) | 28.9 | 30.4 | 63.7 | 65.9 | 62.8 | 44.1 | 43.2 | 79.3 | 90.4 | 70.4 | 78.9 |
| MTL(256128) | 31.2 | 32.1 | 65.0 | 66.2 | 63.9 | 46.3 | 47.6 | 77.1 | 91.2 | 72.4 | 81.9 | |
| PASS [26] | STL(256128) | 32.2 | 35.8 | 64.3 | 64.5 | 66.3 | 46.2 | 46.5 | 80.3 | 92.1 | 72.2 | 83.3 |
| MTL(256128) | 35.3 | 37.3 | 66.4 | 66.1 | 67.1 | 46.7 | 47.7 | 78.4 | 93.3 | 72.5 | 84.7 | |
| STL(384192) | 32.5 | 36.2 | 64.6 | 64.8 | 67.1 | 46.7 | 46.6 | 80.2 | 92.7 | 73.2 | 84.0 | |
| MTL(384192) | 35.7 | 37.9 | 67.7 | 66.6 | 67.5 | 46.9 | 48.1 | 78.8 | 93.7 | 74.5 | 86.1 | |
Ablation on different pre-trained models To investigate the impact of different pre-trained models on model performance, we provide more results about IRM and IRM++ on the publicly released pre-trained models i.e., DeiT [87], ALBEF [74], HAP [88], and PASS [26]. We validate IRM based on the task-specific evaluation setting as shown in Tab. IX and test IRM++ on the task-free evaluation setting as shown in Tab. X. We can see two conclusions from the results. First, we can see models pre-trained on human-centric images, i.e., HAP [88] and PASS [26], can naturally increase the performance of person-retrieval tasks. Second, although ALBEF [74] exhibits lower performance on Trad-ReID and other image-based ReID tasks, it achieves the highest performance on the T2I-ReID task due to its vision-language pretraining knowledge. Additionally, we validate the effectiveness of IRM on 384192 image resolution using the PASS pre-trained model. The results show a further performance improvement compared to the 256128 results, which indicates that increasing the image size can contribute to better retrieval results.
@@ V-F Visualization
Visualization of mAP. In this subsection, we visually validate the improvement of the proposed evaluation metric over the traditional mAP metric. We visualize the indexing results on the CTCC-ReID, LI-ReID, and T2I-ReID tasks, setting the threshold to 0.75. The visualization demonstrates that mAP provides a more accurate measure of both identity correctness and retrieval description consistency in the retrieval results. The green box indicates correctly indexed positive samples, the red box indicates incorrectly indexed negative samples and the yellow box indicates indexed results with correct IDs but with instruction similarity below the set threshold, which are then relabeled as negative samples. For instance, in Fig 12(a) for person 1, the 4th image, and in Fig 12(b) for person 1, the 3rd image, although the model successfully indexed images with the same ID as the query image, it did not find the person wearing the blue jacket as required by the query instruction. In these scenarios, the proposed mAP evaluation metric can successfully filter out retrieval results that do not match the query instruction description. In addition to assessing whether clothing features match, this evaluation metric can also validate the ability of the ReID model to recognize fine-grained features such as body posture and accessories. For example, in Fig 12(b) for person 2, the 5th image (body orientation does not match the retrieval target), and in Fig 12(c) for person 2, the 5th image (unable to discern the status of the carried accessory). Setting the threshold for mAP is a crucial consideration, as seen in Fig 12(c) for person 1, the 4th and 5th images, where an excessively large threshold may classify some true positive samples as negative samples. Choosing an appropriate threshold is a topic worthy of future research.

### v2

@@ Abstract
Recently, person re-identification (ReID) has witnessed fast development due to its broad practical applications and proposed various settings, e.g., traditional ReID, clothes-changing ReID, and visible-infrared ReID. However, current studies primarily focus on single specific tasks, which limits model applicability in real-world scenarios. This paper aims to address this issue by introducing a novel instruct-ReID task that unifies 6 existing ReID tasks in one model and retrieves images based on provided visual or textual instructions. Instruct-ReID is the first exploration of a general ReID setting, where 6 existing ReID tasks can be viewed as special cases by assigning different instructions. To facilitate research in this new instruct-ReID task, we propose a large-scale OmniReID++ benchmark equipped with diverse data and comprehensive evaluation methods, e.g., task-specific and task-free evaluation settings. In the task-specific evaluation setting, gallery sets are categorized according to specific ReID tasks. We propose a novel baseline model, IRM, with an adaptive triplet loss to handle various retrieval tasks within a unified framework. For task-free evaluation setting, where target person images are retrieved from task-agnostic gallery sets, we further propose a new method called IRM++ with novel memory bank-assisted learning. Extensive evaluations of IRM and IRM++ on OmniReID++ benchmark demonstrate the superiority of our proposed methods, achieving state-of-the-art performance on 10 test sets. The datasets, the model, and the code will be available at https://github.com/hwz-zju/Instruct-ReID.
@@
Task-free evaluation setting. When evaluating the generalization ability of the model across multiple retrieval tasks jointly, task-specific instructions for each gallery image are discarded. Therefore, we propose a novel task-free evaluation setting as an extension of our previous conference paper. As shown in Fig. 2(b), the test set maintains a consistent gallery image input for all tasks, with each image in the gallery set being task-agnostic. When extracting gallery data features, only the image features are extracted for various ReID tasks. The gallery features can be fixed and extracted only once in applications, thereby avoiding re-extraction across various application scenarios in the task-specific evaluation setting, which saves computational resources. -
@@
To address the instruct-ReID task, we present an Instruct-ReID Model (IRM) in our previous conference version that can, for the first time, handle the major 6 ReID tasks listed in Fig. 1 (a) with a dual-branch framework. Unlike the typical triplet loss [25] that only defines positive/negative pairs by identities, our IRM incorporates a novel adaptive triplet loss to learn a metric space that preserves identity and instruction similarities. Specifically, we design an adaptive margin between two query-instruction pairs based on instruction similarities to pull features with similar instructions close and push features with different instructions apart. Although the previous IRM has achieved significant success in unifying the multiple ReID tasks within one framework, it can not be applied in the newly proposed task-free evaluation setting for the following two reasons. First, the gallery features of various ReID tasks are extracted from different modules of IRM, e.g., Trad-ReID from the output of the attention module and T2I-ReID from the image encoder. However, in the task-free evaluation setting, all images in the gallery set are task-agnostic, making it impractical to extract retrieval features from different model parts based on the task type of gallery images. Second, IRM extracts gallery images and specific instruction features for retrieval comparison, while these task-specific instructions are unavailable in the task-free evaluation setting. To address the challenges, we introduce a new method called IRM++, which employs simple and unified gallery features for various ReID tasks. Concretely, unlike IRM, IRM++ obtains the gallery features by only extracting the image features from diverse tasks using an image encoder, without introducing task-specific multimodality instruction inputs. Furthermore, IRM++ introduces two memory banks in a contrastive learning manner to effectively supervise the model learning. Unlike triplet loss-based methods that include only one negative sample per triplet, memory bank contrastive learning offers multiple negative samples for each iteration, enhancing the discriminative representation learning of the data.
@@ II-B Loss Function in ReID
Mulitple loss function have been extensively studied during the development of the ReID research [28]. Depending on the structure of the method design, there are three typical types of loss function: (1) Identity loss proposed in [30, 31, 32, 33, 34, 35] treats each identity as a distinct class, thus turning the training phase of ReID as a classification process. Identity loss using cross-entropy loss to compute the difference between the labels and the predicted probability. It’s convenient for training but easy to overfit when the labels are over-confident annotated [36]. (2) Contrastive loss [37, 38, 39, 40, 41, 42] has been widely used in diverse ReID scenarios. It utilizes Euclidean distance to compute the pairwise distance of two input sample features. Contrastive loss encourages features of the same identities to come closer together while pushing features from different identities farther apart. The Contrastive loss can alleviate the impact of wrong classifications and noise during the clustering stage. Although contrastive loss can be sensitive to the choice of positive and negative samples during training, it faces the challenge of lacking a reference distance between samples in ReID task applications. (3) Triplet loss [43, 44, 45, 46] is also an extensively explored feature clustering method. It uses a retrieval ranking perspective to address the training process of the ReID model [28]. Triplet loss takes one anchor sample, one positive sample, and one negative sample as input; based on the assumption that the distance between the anchor and the positive sample should be smaller than the distance between the anchor and the negative sample by a predefined margin, triplet loss optimizes the distance within the triplet samples during the training phase.
@@ III OmniReID++ Benchmark
| dataset | image | ID | scene | labeled | camera view | crop size | resolution | domain |
| MSMT17 [3] | 30,248 | 1,041 | 15 | yes | fixed | vary | fixed | indoor/outdoor |
| Market1501 [23] | 12,936 | 751 | 6 | yes | fixed | 12864 | fixed | outdoor |
| PRCC [24] | 17,896 | 150 | 3 | yes | fixed | vary | fixed | indoor |
| COCAS+ Real1 [12] | 34,469 | 2,800 | 30 | yes | fixed | 256128 | fixed | indoor/outdoor |
| LLCM [64] | 30,921 | 713 | 9 | yes | fixed | vary | fixed | indoor/outdoor |
| LaST [65] | 71,248 | 5,000 | - | yes | dynamic | vary | dynamic | indoor/outdoor |
| MALS [63] | 1,510,330 | 1,510,330 | - | yes | dynamic | 531208 | fixed | synthesis |
| LUPerson-T [66] | 957,606 | - | - | pseudo | dynamic | vary | dynamic | indoor/outdoor |
| OmniReID++ | 5,072,218 | 333,825 | 21,697 | yes | dynamic | vary | dynamic | indoor/outdoor/synthetic |
@@ V-C Task-specific Evaluation Setting Results
V-D Task-free Evaluation Setting Results
Single Test Set Evaluation. We evaluate the performance of IRM++ on each single test set, which is a subset of the task-free evaluation setting. As shown in Tab. VI, for clothes template based clothes-changing ReID (CTCC-ReID), language-instructed ReID (LI-ReID), clothes-changing ReID (CC-ReID), visible-infrared ReID (VI-ReID), text-to-image ReID (T2I-ReID), and traditional ReID (Trad-ReID), under Multi-task learning (MTL) training, our IRM++ achieves state-of-the-art (SOTA) results, surpassing the best-performing methods by +0.3% to +22.4% mAP. Experimental results demonstrate that IRM++ effectively extracts generic gallery features and compares them with features extracted from query images and instructions for accurate retrieval. Besides, we conduct a baseline method IRM* by employing a triplet loss as the supervision function. For CTCC-ReID and LI-ReID tasks, IRM++ surpasses the baseline method by +10.0% and +9.0% mAP (STL), +7.1% and +5.8% mAP (MTL), which indicates that the memory bank assisted contrastive learning utilized in IRM++ further enhances the model performance compared to optimization with triplet loss on these two ReID tasks. We analyze that the memory bank-assisted contrastive loss employs multiple negative samples for each iteration, providing stronger supervision during training compared to triplet-based optimization methods, which only utilize one negative sample per iteration. For CC-ReID, VI-ReID, and Trad-ReID, our IRM++ and baseline method achieve similar results, with experimental variations within 0.3% mAP under Single-task learning (STL) training.
Joint Test Set Evaluation. As shown in Tab. VII, we unify all testing set data into a joint set and provide benchmark testing results on the proposed dataset. We also perform inference with existing state-of-the-art methods on this testing set for comparison. Since DualBCT Net [12], RaSa [18], PASS [26], CAL [7] and DEEN [64] cannot receive multimodal instruction inputs, e.g., clothes template images and description texts, and the instruction similarity is unavailable, we only measure the traditional mAP metric on the Joint Test dataset. We also provide the inference performance of the baseline method IRM* on the newly proposed testing set. Under the traditional mAP evaluation metric, our IRM++ achieves the best result of 65.3% mAP, surpassing the baseline method and existing methods by 9.1% mAP and 22.1% mAP, respectively. The results demonstrate that IRM++ exhibits outstanding performance in the task-free evaluation setting. Based on instruction similarity and our proposed mAP evaluation metric, we provide evaluation results at three different thresholds of 0.25, 0.50, and 0.75. Compared to the baseline method, IRM++ achieves better results of +8.9% (mAP25), +6.7% (mAP50), and +6.3% (mAP75). We observe that under the mAP evaluation metric, the test results of our proposed IRM++ show a decreasing trend as the threshold increases. This indicates that with increasing thresholds, more gallery samples with the same identity as the query image but not matching the instruction attributes are corrected back to negative samples, aligning with our proposed mAP evaluation metric. The selection of thresholds for different models and scenarios is an area worthy of future research.
@@ V-F Visualization
Visualization of mAP. In this subsection, we visually validate the improvement of the proposed evaluation metric over the traditional mAP metric. We visualize the indexing results on the CTCC-ReID, LI-ReID, and T2I-ReID tasks, setting the threshold to 0.75. The visualization demonstrates that mAP provides a more accurate measure of both identity correctness and retrieval description consistency in the retrieval results. The green box indicates correctly indexed positive samples, the red box indicates incorrectly indexed negative samples and the yellow box indicates indexed results with correct IDs but with instruction similarity below the set threshold, which are then relabeled as negative samples. For instance, in Fig. 12 (a) for person 1, the 4th image, and in Fig. 12 (b) for person 1, the 3rd image, although the model successfully indexed images with the same ID as the query image, it did not find the person wearing the blue jacket as required by the query instruction. In these scenarios, the proposed mAP evaluation metric can successfully filter out retrieval results that do not match the query instruction description. In addition to assessing whether clothing features match, this evaluation metric can also validate the ability of the ReID model to recognize fine-grained features such as body posture and accessories. For example, in Fig. 12 (b) for person 2, the 5th image (body orientation does not match the retrieval target), and in Fig. 12 (c) for person 2, the 5th image (unable to discern the status of the carried accessory). Setting the threshold for mAP is a crucial consideration, as seen in Fig. 12 (c) for person 1, the 4th and 5th images, where an excessively large threshold may classify some true positive samples as negative samples. Choosing an appropriate threshold is a topic worthy of future research.